OpenStudy (anonymous):

(x^3 -3x)^6. find the term in x^12

OpenStudy (anonymous):

In the expansion of (x+a)^n \[general~ term ~T_{r+1}=C_{r}^n{x}^{n-r}a^r\] calculate general term.

OpenStudy (anonymous):

write x^3 in place of x and -3x in place of a. n=6

OpenStudy (anonymous):

\[T _{r+1}=C_{r}^{6}\left( x^3 \right)^{6-r}\left( -3x \right)^r =C _{r}^{6}x ^{18-3r}\left( -1 \right)^r3^rx^r\] \[=\left( -1 \right)^r3^rC _{r}^{6}x ^{18-3r+r}=\left( -1 \right)^r3^rC _{r}^{6}x ^{18-2r}\] Put 18-2r=12 2r=18-12=6 r=3 \[C _{3}^{6}=\frac{ 6\times5\times4 }{ 3\times2\times1 }=20\] simplify and find the required term.